Find the sum of 1,1/5,1/25


1
Expert's answer
2021-09-17T03:44:39-0400

"1+\frac{1}{5}+\frac{1}{25}=\frac{25}{25}+\frac{5}{25}+\frac{1}{25}=\frac{25+5+1}{25}=\frac{31}{25}=1.24"


Answer: "\frac{31}{25}" , or 1.24 in decimal form